# SubRegWeigh: Effective and Efficient Annotation Weighing with Subword Regularization

### v1

@@ Abstract
Many datasets of natural language processing (NLP) sometimes include annotation errors. Researchers have attempted to develop methods to reduce the adverse effect of errors in datasets automatically. However, an existing method is time-consuming because it requires many trained models to detect errors. We propose a novel method to reduce the time of error detection. Specifically, we use a tokenization technique called subword regularization to create pseudo-multiple models which are used to detect errors. Our proposed method, SubRegWeigh, can perform annotation weighting four to five times faster than the existing method. Additionally, SubRegWeigh improved performance in both document classification and named entity recognition tasks. In experiments with pseudo-incorrect labels, pseudo-incorrect labels were adequately detected.
@@ 1 Introduction
Many datasets of natural language processing (NLP) consist of raw texts and annotation labels. Various NLP tasks exploit the pair of the raw text and the annotation label for training and evaluating models. For example of named entity recognition (NER), which is applied to various practical technologies such as location detection Inkpen et al. (2017) and anonymization Mamede et al. (2016), some parts of the text are annotated as named entities (e.g., location names or personal names). And then, a model is trained to extract these entities from the raw text. To achieve higher performance in NLP tasks, the models should be trained or fine-tuned with a sophisticated training dataset without annotation errors.
@@
However, in fact, some existing popular datasets such as CoNLL-2003 contain annotation errors Wang et al. (2019). When the annotation errors are included in datasets, the performance is degraded by training models from incorrect training datasets and/or models are incorrectly evaluated by errors of test datasets.
In the case of NER, these annotation errors are typically related to ambiguous categories of named entities. For example, the string "America" can refer to either a geographical location or an organization, depending on the context. Additionally, when followed by "Online", the correct entity span should be "America Online", which is the name of an organization, rather than just "America".
@@
It has been reported that CoNLL-2003, the well-known NER dataset, contains annotation errors in both training and test splits Wang et al. (2019); Reiss et al. (2020). Most recently, researchers have also explored automated annotation using generative AI Goel et al. (2023); Bogdanov et al. (2024) and collaborative annotation involving humans and AI Naraki et al. (2024). Even with these methods, it is difficult to avoid annotation errors. Therefore sophisticated weighing methods to automatically detect the annotation errors and reduce their negative effect have been expected.
@@
Such a method to weigh annotation errors is recently studied in the NER field. Wang et al. (2019) proposed CrossWeigh which is the method for detecting annotation errors in the dataset and adjusting their learning priority by weighting loss values so that the training is not affected by such annotation errors. However, there are shortcomings in its computational efficiency, especially in the recent NLP trends with the pre-trained large language models. We consider that the more efficient methods of annotation weighing can speed up the development of NLP. In addition, reducing the computational cost contributes to Green AI Schwartz et al. (2019). Furthermore, we cannot utilize CrossWeigh for NLP tasks other than NER because its architecture is specialized to NER. Therefore, we are required to develop the new weighing methods that can be widely used for various NLP tasks.
@@
Figure 1: The overview of the three steps of SubRegWeigh with the number of tokenization candidates in the inference $K=3$: 1) the training step of the scouting model, 2) the inference step to examine the appropriateness of the training label, and 3) the weighing step to calculate the weights for each training sample according to the appropriateness of their labels. The calculated weights are used to train the final model.
@@
In this study, we propose an efficient and effective method of annotation weighing, SubRegWeigh. SubRegWeigh utilizes the idea of subword regularization based on multiple segmentation results, to detect the annotation errors and weight the loss values to degrade the effect of training samples with errors. Figure 1 overviews the proposed method. Inspired by Takase et al. (2022), we obtain the multiple outputs from a single trained model by inputting multiple variations of tokenization candidates. In other words, SubRegWeigh measures annotation reliability by observing the consistency of multiple outputs generated with subword regularization. SubRegWeigh can be regarded as a method to create pseudo-multiple models for annotation evaluation. This allows for faster weighing of annotation errors by using a single model instead of multiple models required by CrossWeigh.
@@
We conducted experiments on two NLP tasks: NER and text classification. The experimental results of the CoNLL-2003 demonstrate that SubRegWeigh can perform the annotation weighing four to five times faster than CrossWeigh. Furthermore, SubRegWeigh contributes to the performance improvement of the trained model compared to the case using CrossWeigh in some datasets of both text classification and NER tasks. Especially, our proposed method achieved SoTA on CoNLL-CW, a test dataset constructed by manually correcting annotation errors in the CoNLL-2003 test split. Experiments with pseudo-incorrect labels also showed that the proposed method can recognize annotation errors well.
@@ 2 Related Work
2.1 Annotation Error Detecting
Our work is in the line of CrossWeigh Wang et al. (2019). CrossWeigh detects annotation errors by performing $K$-fold cross-validation on the training dataset $T$ times and weighs the loss value according to the number of wrong predictions, which degrades the effect of training samples with annotation errors in the NER task. This method requires large computational costs because it requires training $K\times T$ models for detecting annotation errors. Our work is different from CrossWeigh in that we use multiple tokenization candidates to obtain various outputs from one model instead of $K$-fold cross-validation. Compared to CrossWeigh, the proposed method can remarkably reduce the time for detecting annotation errors because we train the model only once and use the output from pseudo-models by inputting various tokenization candidates. Furthermore, our method can use the entire training dataset to make the model because we do not employ $K$-fold cross-validation, which contributes to the more accurate detection of annotation errors.
In addition to methods that pre-assign annotation correctness weights to the training data set, methods have been studied to correct labels in partially annotated training data using a small set of data known to have correct annotations Yang et al. (2018); Mayhew et al. (2019); Ding et al. (2023). In other lines, some methods are developed to infer the correctness of the labels during training. For instance, in the method introduced in Zhou and Chen (2021), multiple models are trained simultaneously with different initial values, and the KL divergence of outputs for the same sentence is added to the loss function. This method can be applied in conjunction with the weighting of training datasets used by CrossWeigh and our method.
For detecting annotation errors in datasets beyond NER, annotation correction methods with distant supervision Mintz et al. (2009) have been studied in the relation extraction tasks Huang and Du (2019); Qin et al. (2018). However, these methods require an additional clean dataset or are not always applicable. Furthermore, methods for correcting POS tagging tasks have also been studied Nakagawa and Matsumoto (2002); Helgadóttir et al. (2014), but these do not apply to tasks other than POS tagging. Maini et al. (2023) introduced a method to analyze the memorization of the model, which can be used to detect training samples with ambiguous labels.
@@ 2.2 Subword Regularization
The proposed method exploits the technique of tokenization for the efficient detection of annotation errors. Tokenization Song et al. (2021); Sennrich et al. (2016); Kudo (2018) is widely used in recent natural language processing. This technique reduces the number of unknown words and low-frequency words and limits the number of dictionaries by treating words as a sequence of finer subwords. However, because tokenization is deterministic processed on training data, the subword sequences that appear during inference may not always match the sequences learned during training.
Subword regularization Kudo (2018); Provilkov et al. (2020); Hiraoka (2022) alleviates this problem of deterministic tokenization. This technique generates multiple tokenization candidates for a single text. While subword regularization is typically used during the training step to learn from various tokenizations, some studies use it during inference. Takase et al. (2022) uses subword regularization during inference to ensemble the various outputs of a single model by inputting some different tokenization candidates to the model. Wang et al. (2021a) proposed a method to reduce the effect of tokenization difference. However, there has been no research on using subword regularization during inference to reduce noise in the training data.
Subword regularization using a unigram language model Kudo (2018) is a method to make the model more robust by sampling tokenization candidates using the unigram probability of tokens. For the cases where the model does not employ a tokenizer with the unigram language model like Byte Pair Encoding (BPE) Sennrich et al. (2016) used in RoBERTa Liu et al. (2019) or WordPiece Song et al. (2021) used in BERT Devlin et al. (2019), we can use BPE-Dropout Provilkov et al. (2020) and MaxMatch-Dropout Hiraoka (2022) for the subword regularization, respectively. BPE-Dropout randomly avoids merges during the combination process to obtain multiple subword sequences. MaxMatch-Dropout randomly rejects the longest match candidates for each word to obtain multiple subword sequences.
@@ 3 Proposed Method: SubRegWeigh
The purpose of the proposed method, SubRegWeigh, is to detect the incorrectly annotated samples in the training dataset and reduce the effect of such a sample with annotation errors during the training of the final model by weighting the loss values. The process of SubRegWeigh is composed of the following three steps:
@@
    Training a scouting model in a usual manner of NLP tasks with the training data. Deterministic tokenization is used for this training (§3.1).
@@
    Inferring the labels for the inputs in the training dataset with the trained scouting model. Herein, we tokenize the input texts into $K$ different tokenization candidates. Therefore, we can collect $K$ inferred outputs for each tokenization candidates of a single input text. (§3.2)
@@
    Calculating the weights for the training sample according to the number of correctly answered inputs for each tokenization candidate (§3.3).
This process is overviewed in Figure 1. The first and second steps detect the incorrect annotations by scouting the dataset before the final model training. Then, SubRegWeigh weighs the correctness of annotation labels and assigns the weight for each training sample in the third step. After the above steps, we can train the final model with the original training data and the calculated weights for each training sample (§3.4). Compared to CrossWeigh, SubRegWeigh can be used for various NLP tasks beyond NER because the entire steps are not specialized to NER. The following section explains each step in detail.
@@ 3.1 Training of the Scouting Model
Inspired by CossWeigh Wang et al. (2019), we first create the scouting model $M$ (see the left top of Figure 1). This model is used to scout the correctness of the annotated label in the training dataset in the next step. Let $(X,Y)$ be a pair of the text and the label corresponding to a single training sample. When the text is composed of $I$ words, $X=x_{1},...,x_{I}$. $Y$ is a single value in the text classification (e.g., positive and negative labels for sentiment classification) or a sequence of labels $Y=y_{1},...,y_{I}$ in sequential labeling tasks such as NER.
In this step, we train the scouting model $M$ in the usual manner of training or fine-tuning for general NLP datasets. The single input text $X$ is tokenized into a sequence of subwords $X^{\prime}=s_{1},...,s_{J}$ that is composed of $J$ subwords [^1] The numbers of words $I$ and subwords $J$ are not necessarily matched depending on the tokenization methods. In the case of sequential labeling tasks with $I\neq J$, we aligned the subwords and labels following the existing literature Liu et al. (2019); Yamada et al. (2020).. After tokenizing all texts in the training dataset, we train the scouting model $M$ with the tokenized texts and labels. Note that we use the deterministic tokenization here (e.g., the default encoding of the BPE tokenizer or the $1$-best tokenization with the Viterbi algorithm for the unigram language model-based tokenizer). In other words, we do not use subword regularization (i.e., stochastic tokenization) for the training of the scouting model $M$.
@@ 3.2 Inference with the scouting Model
In the second step, we use the trained scouting model $M$ to scout the correctness of the annotated label in the training dataset (see the right part of Figure 1). We assume that the trained scouting model $M$ can correctly predict the labels even for the differently tokenized input if the label is not wrongly annotated. Specifically, the input text $X$ is tokenized into $K$ different tokenization candidates $\{X^{\prime}_{1},...,X^{\prime}_{k},...,X^{\prime}_{K}\}$. Then each tokenization candidate $X^{\prime}_{k}$ is fed into the trained model $M$ and we obtain the corresponding prediction $\hat{Y}_{k}$ [^2] As well as $Y$, $\hat{Y}_{k}$ can be single or multiple labels depending on target tasks. After the inference, we evaluate the output $\hat{Y}_{k}$ with the gold annotation $Y$. This process can be regarded as the “mistake-reweighing” process of CrossWeigh with $K$ pseudo-models, which is realized with the $K$ different tokenization candidates for the single input text and the single model.
@@
We consider it important to make a variety of the $K$ tokenization candidates to weigh the correctness of the annotation from diverse aspects. To avoid using similar subword sequences as the inputs for the model $M$, we select $K$ much different candidates from a large number $N$ of tokenization candidates. We propose three options of the way to select such $K$ tokenization candidates as follows.
@@
As the most straightforward way, we randomly sample $N=K$ tokenization candidates using subword regularization and use them as inputs for the model $M$. In other words, we do not select the tokenization candidates from the large number of candidates.
@@
We select $K$ tokenization candidates using the cosine similarity calculated with TF-IDF vectors. TF-IDF was calculated with each subword as a term, each subword sequence as a document, and the entire candidate set as a corpus. First, we convert $N(>K$) tokenization candidates into TF-IDF vectors and select the first candidate with the smallest cosine similarity against the deterministic tokenization $X^{\prime}$. Subsequently, we select the candidate that is not similar both to $X^{\prime}$ and already selected candidates. We repeatedly select the candidates until the number of candidates reaches $K$.
@@
We select $K$ tokenization candidates using the $K$-means clustering. We apply the clustering to $N(>K)$ candidates vectorized with TF-IDF. Then we select $K$ candidates which are located at the nearest point to the centroids of each cluster.
@@ 3.3 Weighting for Training Samples
To reduce the effect of training samples with annotation errors, we calculate weights for each sample depending on the results of the inference step (see the left bottom of Figure 1). Concretely, we calculate the weight $w$ corresponding to the training sample $(X,Y)$ as the follows:
@@
where $\hat{Y}_{k}$ is the output of the scouting model $M$ when inputting the tokenization candidate $X^{\prime}_{k}$. If all predictions were different from the original label $Y$, the weights would be 0 and the data would not be used for training. To avoid this, we use a pre-defined minimum weight $w_{\mathrm{min}}$. We use this minimum weight because Wang et al. (2019) reported that it is better to give a small weight to samples with annotation errors than not to use them.

### v2

@@ Abstract
NLP datasets may still contain annotation errors, even when they are manually annotated. Researchers have attempted to develop methods to automatically reduce the adverse effect of errors in datasets. However, existing methods are time-consuming because they require many trained models to detect errors. This paper proposes a time-saving method that utilizes a tokenization technique called subword regularization to simulate multiple error detection models for detecting errors. Our proposed method, SubRegWeigh, can perform annotation weighting four to five times faster than the existing method. Additionally, SubRegWeigh improved performance in document classification and named entity recognition tasks. In experiments with pseudo-incorrect labels, SubRegWeigh clearly identifies pseudo-incorrect labels as annotation errors. Our code is available at https://github.com/4ldk/SubRegWeigh.
@@ 1 Introduction
NLP datasets usually consist of raw texts and annotation labels. Various tasks exploit the pairs of texts and labels for training and evaluating models. To achieve higher performance in NLP tasks, the models should be trained or fine-tuned with a sophisticated training dataset without annotation errors.
@@
However, some popular datasets such as CoNLL-2003 contain annotation errors Wang et al. (2019); Reiss et al. (2020). When datasets include errors, the performance is degraded by training models from incorrect training datasets, and models are incorrectly evaluated by errors of test datasets.
@@
Recent studies have explored automated annotation using generative AI Goel et al. (2023); Bogdanov et al. (2024) and the collaboration of humans and AI Naraki et al. (2024). Even with these methods, it is difficult to avoid annotation errors. Therefore sophisticated weighing methods to automatically detect the annotation errors and reduce their negative effect have been expected.
@@
Such a method to weigh annotation errors is recently studied in the NER field. Wang et al. (2019) proposed CrossWeigh, which detects annotation errors in the dataset and adjusts their learning priority by weighting loss values so that the training is not affected by such annotation errors. However, there are shortcomings in its computational efficiency, especially in the recent NLP trends with the pre-trained large language models. Reducing the computational cost contributes to not only speeding up the development of NLP but also Green AI Schwartz et al. (2019). Furthermore, we cannot utilize CrossWeigh for NLP tasks other than NER because it is specially designed for NER. Therefore, we are required to develop the new weighing methods that can be widely used for various NLP tasks.
@@
Figure 1: The overview of the three steps of SubRegWeigh with the number of tokenization candidates in the inference $K=3$ for the classification task whether the input sentence is positive or negative: 1) the training step of the scouting model, 2) the inference step to examine the appropriateness of the training label, and 3) the weighting step to calculate the weight of each training sample according to the appropriateness of their labels. Check marks indicate inferences that are the same as the original label and cross marks indicate inferences that differ from the original label. The calculated weights are used to train the final model.
@@
In this study, we propose an efficient and effective method of annotation weighing, SubRegWeigh. Figure 1 shows an overview of the proposed method. SubRegWeigh evaluates the reliability of annotation for a single sample by feeding various input styles to a model and comparing its outputs. Inspired by Takase et al. (2022), we obtain such various inputs with subword regularization, which yields various tokenization candidates. In other words, SubRegWeigh measures annotation reliability to detect errors by observing the consistency of multiple outputs generated with subword regularization. SubRegWeigh can also be regarded as a method to simulate multiple models for annotation evaluation, which allows for faster weighing of annotation errors by using a single model instead of multiple models required by CrossWeigh. Because the weighing processing of SubRegWeigh is not ad hoc for NER, it is applicable to wide NLP tasks.
@@
We conducted experiments on two NLP tasks: NER and text classification. The experimental results of the CoNLL-2003 demonstrate that SubRegWeigh can perform the annotation weighing four to five times faster than CrossWeigh. Furthermore, SubRegWeigh contributes to the performance improvement compared to the case using CrossWeigh in some datasets. Especially, our proposed method achieved SoTA on CoNLL-CW, a test dataset constructed by manually correcting annotation errors in the CoNLL-2003 test split. Experiments with pseudo-incorrect labels also showed that the proposed method can recognize annotation errors well.
@@ 2 Related Work
2.1 Annotation Error Detection
Our work is in the line of CrossWeigh Wang et al. (2019), which is an annotation weighing method for NER. CrossWeigh detects annotation errors by performing $K$-fold cross-validation on the training dataset $T$ times and weights the loss value according to the number of wrong predictions, which degrades the effect of training samples with annotation errors. This method requires large computational costs because it trains $K\times T$ models for detecting annotation errors. Our work is different from CrossWeigh in that we use multiple tokenization candidates to obtain various outputs from one model. This can remarkably reduce the time for detecting annotation errors because we train the model only once and use the output from pseudo-models by inputting various tokenization candidates. Furthermore, freed from $K$-fold cross-validation, our method can use the entire training dataset to make the model, which contributes to the more accurate detection of annotation errors.
For the case where we can prepare a small dataset with absolutely correct annotations, some methods have been studied to correct labels directly Yang et al. (2018); Mayhew et al. (2019); Ding et al. (2023). In other lines, Zhou and Chen (2021) proposed a method to infer the correctness of the labels during training, which can be applied in conjunction with our method.
For detecting annotation errors in datasets beyond NER, annotation correction methods with distant supervision Mintz et al. (2009) have been studied in the relation extraction tasks Huang and Du (2019); Qin et al. (2018). However, these methods require an additional clean dataset or are not always applicable. Furthermore, methods for correcting POS tagging tasks have also been studied Nakagawa and Matsumoto (2002); Helgadóttir et al. (2014), but these do not apply to tasks other than POS tagging. The analyzing method of the models’ memorization Maini et al. (2023) could be used to detect training samples with ambiguous labels.
@@ 2.2 Subword Regularization
The proposed method exploits the technique of tokenization for efficient annotation error detection. Tokenization Song et al. (2021); Sennrich et al. (2016); Kudo (2018) is widely used in recent NLP. This technique reduces the number of unknown and low-frequency words and limits the number of dictionaries by treating words as a sequence of finer subwords. However, because tokenization is deterministically performed on training data, the subword sequences of inference data may not always match the sequences learned during training.
Subword regularization Kudo (2018); Provilkov et al. (2020); Hiraoka (2022) alleviates this problem of deterministic tokenization. This technique generates multiple tokenization candidates for a single text. While subword regularization is typically used during the training step to learn from various tokenizations, some studies use it during inference. Takase et al. (2022) uses subword regularization during inference to ensemble the various outputs of a single model by inputting some different tokenization candidates to the model. Wang et al. (2021a) proposed a method to reduce the effect of tokenization difference. However, there has been no research using subword regularization during inference to reduce noise in the training data.
Subword regularization using a unigram language model Kudo (2018) is a method to sample tokenization candidates using the unigram probability of tokens. For the cases where the model does not employ a tokenizer with the unigram language model like Byte Pair Encoding (BPE) Sennrich et al. (2016) used in RoBERTa Liu et al. (2019) or WordPiece Song et al. (2021) used in BERT Devlin et al. (2019), we can use BPE-Dropout Provilkov et al. (2020) and MaxMatch-Dropout Hiraoka (2022) for the subword regularization, respectively. To obtain various tokenization candidates, BPE-Dropout randomly avoids merges during the combination process and MaxMatch-Dropout randomly rejects the longest match candidates for each word.
@@ 3 Proposed Method: SubRegWeigh
The purpose of the proposed method, SubRegWeigh, is to detect the incorrectly annotated samples in the training dataset and reduce the effect of such a sample with annotation errors during the training of the final model by weighting the loss values. The process of SubRegWeigh is composed of the following three steps (Figure 1):
@@
    Training a scouting model in a usual manner of NLP tasks with the training data using deterministic tokenization (§3.1).
@@
    Inferring the labels for the training samples with the trained scouting model. Herein, we tokenize the input texts into $K$ different tokenization candidates. Therefore, we can collect $K$ inferred outputs for each tokenization candidate of a single input text (§3.2).
@@
    Calculating weights for training samples according to the number of correctly answered inputs for each tokenization candidate (§3.3).
The first and second steps detect the annotation errors by scouting the dataset. Then, SubRegWeigh weighs the correctness of annotation labels and assigns the weight for each training sample in the third step. After the above steps, we can train the final model with the original training data and the calculated weights for each training sample (§3.4). Compared to CrossWeigh, SubRegWeigh can be used for various NLP tasks beyond NER because the entire steps are not specialized to NER. The following section explains each step in detail.
@@ 3.1 Training of the Scouting Model
Inspired by CossWeigh Wang et al. (2019), we first create the scouting model $M$ (see the left top of Figure 1). This model is used to scout the correctness of the annotation in the training dataset in the next step. Let $(X,Y)$ be a pair of the text and the label corresponding to a single training sample. When the text is composed of $I$ words, $X=x_{1},...,x_{I}$. $Y$ is a single value in text classification (e.g., positive and negative labels) or a sequence of labels $Y=y_{1},...,y_{I}$ in sequential labeling such as NER.
In this step, we train the scouting model $M$ in the usual manner of training or fine-tuning for general NLP datasets. The single input text $X$ is tokenized into a sequence of subwords $X^{\prime}=s_{1},...,s_{J}$ that is composed of $J$ subwords [^1] The numbers of words $I$ and subwords $J$ are not necessarily matched depending on the tokenization methods. In the case of sequential labeling tasks with $I\neq J$, we aligned the subwords and labels following the existing literature Liu et al. (2019); Yamada et al. (2020).. Then, we train the scouting model $M$ with the tokenized texts and labels. Note that we use the deterministic tokenization here (e.g., the default encoding of the BPE tokenizer or the $1$-best tokenization with the Viterbi algorithm for the unigram language model-based tokenizer). In other words, we do not use subword regularization (i.e., stochastic tokenization) for the training of the scouting model $M$.
@@ 3.2 Inference with the Scouting Model
In the second step, we use the trained scouting model $M$ to scout the correctness of the annotated label in the training dataset (see the left bottom of Figure 1). We assume that the trained scouting model $M$ can correctly predict the labels even for the differently tokenized input if the label is correctly annotated. Concretely, the input text $X$ is tokenized into $K$ different tokenization candidates $\{X^{\prime}_{1},...,X^{\prime}_{k},...,X^{\prime}_{K}\}$. Then each tokenization candidate $X^{\prime}_{k}$ is fed into the trained model $M$ and we obtain the corresponding prediction $\hat{Y}_{k}$ [^2] As well as $Y$, $\hat{Y}_{k}$ can be single or multiple labels depending on target tasks. We then evaluate the output $\hat{Y}_{k}$ with the gold annotation $Y$. This process can be regarded as the “mistake-reweighing” process of CrossWeigh with $K$ pseudo-models, which is realized with the $K$ different tokenization candidates for the single input text and the single model Takase et al. (2022).
@@
We consider it important to make a variety of the $K$ tokenization candidates to weigh the correctness of the annotation from diverse aspects. To avoid using similar subword sequences as the inputs for the model $M$, we select $K$ more different candidates from a large number $N$ of tokenization candidates. We propose three options of the way to select such $K$ tokenization candidates as follows.
@@
We randomly sample $N=K$ different tokenization candidates using subword regularization and use them as inputs for the model $M$. In other words, we do not select the tokenization candidates from the large number of candidates.
@@
We select $K$ tokenization candidates using the cosine similarity calculated with TF-IDF vectors. TF-IDF was calculated with each subword as a term, each subword sequence as a document, and the entire candidate set as a corpus. First, we convert $N(>K$) tokenization candidates into TF-IDF vectors and select the first candidate with the smallest cosine similarity against the deterministic tokenization $X^{\prime}$. Subsequently, we select the candidate that is least similar to both $X^{\prime}$ and already selected candidates. We repeatedly select the candidates until the number of candidates reaches $K$.
@@
We select $K$ tokenization candidates using the $K$-means clustering. We apply the clustering to $N(>K)$ candidates vectorized with TF-IDF. We then choose $K$ candidates at the nearest point to the centroids of each cluster.
@@ 3.3 Weighting for Training Samples
To reduce the effect of training samples with annotation errors, we calculate weights for each sample depending on the results of the inference step (see the right part of Figure 1). Concretely, we calculate the weight $w$ corresponding to the training sample $(X,Y)$ as the follows:
@@
where $\hat{Y}_{k}$ is the output of the scouting model $M$ when inputting the tokenization candidate $X^{\prime}_{k}$. If all predictions were different from the original label $Y$, the weights would be 00 and the data would not be used for training. We use a pre-defined minimum weight $w_{\mathrm{min}}$ to avoid the zero weights, following the suggestion in Wang et al. (2019) finding it better to give a small weight to samples with annotation errors than not to use them.